# Adapting Diarization-Conditioned Whisper for End-to-End Multi-Talker Speech Recognition

### v1

@@ 1 Introduction
Recent advances in multi-talker ASR (MT-ASR) have focused on directly transcribing overlapping speech while handling speaker assignment. Permutation-invariant training [9] addresses the issue of speaker label ambiguity but suffers from factorial complexity, motivating successors such as Heuristic Error Assignment Training [10, 11], which approximates optimal assignments. Serialized output training (SOT) [12] simplifies decoding with speaker change tokens but lacks explicit speaker modeling. To better leverage speaker cues, Diarization-Conditioned Whisper (DiCoW) [13] adapts Whisper with lightweight diarization-conditioned modules, SLIDAR [14] integrates diarization and ASR within a unified model, and Microsoft’s Whisper extension [15] injects speaker change tokens to improve overlap handling.
@@ 2.2 Target-speaker conditioning in DiCoW
where $p^{t}_{\mathcal{S}}$, $p^{t}_{\mathcal{T}}$, $p^{t}_{\mathcal{N}}$ and $p^{t}_{\mathcal{O}}$ correspond to STNO probabilities derived from standalone diarization at time $t$. This convex combination allows the model to adjust its internal representations depending on the speaker context. See [13] for additional details. In this work, we used only the oracle diarization derived from human annotations, i.e. the transformation simplifies to selecting the corresponding class-specific transformation.
@@ 3.1 Datasets
All experiments are conducted using single-channel audio and oracle diarization derived from the available annotations. We train our proposed model using only English multi-talker datasets: NOTSOFAR [24], AMI [17], and LibriMix [16]. Data preparation is handled using Lhotse recipes [25], with minor modifications to ensure that all segments comply with the 30-second input constraint of Whisper.
@@ 4.2 Ablation study on improving speaker label assignment
For reference, DiCoW, that is, a target-speaker model that decodes each speaker independently, achieves 18.4 % cpWER in NOTSOFAR, outperforming our best SA-DiCoW model (20.8 %). However, on LibriMix with 3 speakers, DiCoW obtains 32.1 % cpWER, while our model achieves 17.2 %. This drop in DiCoW’s performance on LibriMix is expected: in fully overlapped conditions, the STNO masks for different speakers become ambiguous, giving the model little guidance on which speaker to transcribe. While this limitation becomes apparent in synthetic mixtures with dense overlap, such extreme conditions are rare in natural conversations.
@@ 4.4 Comparison with prior speaker-attributed ASR systems
Table 4 compares our proposed SA-DiCoW model with existing speaker-attributed ASR systems on AMI-SDM and AMI-IHM-MIX. On AMI-SDM, our method achieves (18.1 %) cpWER, outperforming previous speaker-attributed works such as Cornell et al. (21.1 %) [14] and Li et al. (21.2 %) [15]. This underscores the effectiveness of our approach in handling challenging far-field conversational data. For reference, the original target-speaker DiCoW model achieves the lowest cpWER on both datasets.

### v2

@@ 1 Introduction
Recent advances in multi-talker ASR (MT-ASR) have focused on directly transcribing overlapping speech while handling speaker assignment. Permutation-invariant training [30] addresses the issue of speaker label ambiguity but suffers from factorial complexity, motivating successors such as HEAT [24, 15] or Sortformer [19], which approximates optimal assignments. Serialized output training (SOT) [11] simplifies decoding with speaker change tokens but lacks explicit speaker modeling. To better leverage speaker cues, Diarization-Conditioned Whisper (DiCoW) [21, 20] adapts Whisper with lightweight diarization-conditioned modules, DNCASR [32] and SLIDAR [6] integrate diarization and ASR within a unified model, and Microsoft’s Whisper extension [14] injects speaker change tokens to improve overlap handling.
@@ 2.2 Target-speaker conditioning in DiCoW
where $p^{t}_{\mathcal{S}}$, $p^{t}_{\mathcal{T}}$, $p^{t}_{\mathcal{N}}$ and $p^{t}_{\mathcal{O}}$ correspond to STNO probabilities derived from standalone diarization at time $t$. This convex combination allows the model to adjust its internal representations depending on the speaker context. See [21] for additional details. In this work, we used only the oracle diarization derived from human annotations, i.e. the transformation simplifies to selecting the corresponding class-specific transformation. While these oracle masks result in one-hot vectors, the FDDT mechanism natively supports soft assignments from automatic diarization systems without architectural changes.
@@ 3 Experimental Setup
All experiments are conducted using single-channel audio and oracle diarization derived from the available annotations.
@@ 3.1 Datasets
We train our proposed model using only English multi-talker datasets: NOTSOFAR [27], AMI [2], and LibriMix [8]. Data preparation is handled using Lhotse recipes [31], with minor modifications to ensure that all segments comply with the 30-second input constraint of Whisper.
@@ 4.2 Ablation study on improving speaker label assignment
For reference, DiCoW, that is, a target-speaker model that decodes each speaker independently, achieves 18.4 % cpWER in NOTSOFAR, outperforming our best SA-DiCoW model (20.8 %). However, on LibriMix with 3 speakers, DiCoW obtains 32.1 % cpWER, while our model achieves 17.2 %. This drop in DiCoW’s performance on LibriMix is expected: in fully overlapped conditions, the STNO masks for different speakers become ambiguous, giving the model little guidance on which speaker to transcribe. These results validate the design of SA-DiCoW, demonstrating that access to global conversational context is indeed beneficial, especially in overlap-heavy conditions.
@@ 4.4 Comparison with prior speaker-attributed ASR systems
Table 4 compares our proposed SA-DiCoW model with existing speaker-attributed ASR systems on AMI-SDM and AMI-IHM-MIX with oracle diarization. On AMI-SDM, our method achieves (18.1 %) cpWER, outperforming previous works such as Cornell et al. (21.1 %) [6] and Li et al. (21.2 %) [14]. This underscores the effectiveness of our approach in handling challenging far-field conversational data. For reference, the original target-speaker DiCoW model achieves the lowest cpWER on both datasets.
@@ 5 Conclusion
6 Acknowledgements
The work was supported by Czech Ministry of Culture NAKI III project JARIN (DH23P03OVV010) and by Ministry of Education, Youth and Sports of the Czech Republic (MoE) through the OP JAK project ”Linguistics, Artificial Intelligence and Language and Speech Technologies: from Research to Applications” (ID:CZ.02.01.01/00/23_020/0008518). Computing on IT4I supercomputer was supported by MoE through the e-INFRA CZ (ID:90254).